# From Open Vocabulary to Open World: Teaching Vision Language Models to Detect Novel Objects

### v1

@@ Abstract
Traditional object detection methods operate under the closed-set assumption, where models can only detect a fixed number of objects predefined in the training set. Recent works on open vocabulary object detection (OVD) enable the detection of objects defined by an unbounded vocabulary, which reduces the cost of training models for specific tasks. However, OVD heavily relies on accurate prompts provided by an “oracle”, which limits their use in critical applications such as driving scene perception. OVD models tend to misclassify near-out-of-distribution (NOOD) objects that have similar semantics to known classes, and ignore far-out-of-distribution (FOOD) objects. To address theses limitations, we propose a framework that enables OVD models to operate in open world settings, by identifying and incrementally learning novel objects. To detect FOOD objects, we propose Open World Embedding Learning (OWEL) and introduce the concept of Pseudo Unknown Embedding which infers the location of unknown classes in a continuous semantic space based on the information of known classes. We also propose Multi-Scale Contrastive Anchor Learning (MSCAL), which enables the identification of misclassified unknown objects by promoting the intra-class consistency of object embeddings at different scales. The proposed method achieves state-of-the-art performance in common open world object detection and autonomous driving benchmarks.
@@ 1 Introduction
Recently, the rise of visual-language pre-training has spawned a new area: open vocabulary object detection (OVD). In principle, OVD can detect novel classes defined by an unbounded (open) vocabulary at inference. Naturally, it is able to incrementally learn new classes by adding new prompts. In the real world, however, an open vocabulary may not necessarily be complete and accurate. As shown in Fig. 1, an open vocabulary model such as Yolo World [5] is still a closed-set model at test time, because it is impractical for the expert user, referred to as “oracle”, to define all possible objects that need to be detected. With a limited vocabulary the model tends to misclassify near-out-of-distribution (NOOD) objects that have similar semantics to known classes, and ignore far-out-of-distribution (FOOD) objects. In real-world applications such as autonomous driving, where there are countless object-types that cannot be initially included in the text prompt, OVD will inevitably fail when presented with out-of-distribution objects. The limitation of OVD arises not from the lack of prior knowledge, but from the inability to apply it effectively, a process that requires recognizing the unknown objects as gaps in the model’s representation.
To tackle this challenge, we propose a framework that enables OVD models to operate in open-world settings. To detect FOOD objects and incrementally learn the new classes, we propose Open World Embedding Learning (OWEL). OWEL optimizes parameterized class embeddings rather than fine-tuning the whole model to learn new classes, which inherently reduces catastrophic forgetting. We further introduce the novel concept of Pseudo Unknown Embedding, which constructs a text embedding to detect FOOD objects based on current known classes and generic objectness.
@@
Our method significantly surpasses state-of-the-art (SOTA) performance in U-Recall on the M-OWODB [19] and S-OWODB [14] benchmarks, while maintaining leading performance in other metrics. We further evaluate our method on a novel benchmark based on nuScenes [2], where it also achieves the best results. The contributions of this work are as follows:
  • We propose a framework that enables OVD models to operate in open world settings by identifying unknown objects and incrementally learning new classes.
  • We propose a novel method, Open World Embedding Learning (OWEL), to enable the discovery and incremental learning of new classes without fine-tuning the entire model or requiring exemplars of previously known classes.
@@ 2 Related Works
Closed set object detection has been extensively studied over the past decade [12, 49, 11, 48, 29, 28, 60, 3]. To handle unknown objects and incrementally learn new classes, Joseph et al. [19] first proposed an open world object detector, which extends the Faster R-CNN [49] with contrastive clustering and energy-based unknown-object identification. Subsequent works [14, 35] used contextual information to improve unknown identification and knowledge transfer between known and unknown classes. To effectively detect unknown objects, Zohar et al. [72] proposed Probabilistic Objectness (PROB) to estimate the objectness of different proposals. Wang et al. [64] introduced random proposals in detector training to encourage the unknown discovery and reduce confusion between known and unknown classes. Ma et al. [33] proposed decoupling object localization and classification via cascade decoding. Sun et al. [58] further de-correlates objectness and class information by enforcing orthogonality.
@@
In a recent work [34], the open vocabulary model is used for OWOD via knowledge distillation, however, it still requires exemplar replay-based fine-tuning to incrementally learn new classes.
@@
Out-of-distribution (OOD) detection is a concept closely related to open set learning, and the two terms are sometimes used interchangeably. OOD detection differs slightly from OSR, as OOD samples may be entirely unrelated to the training set, while OSR focuses on rejecting valid but previously unseen classes [10, 51].
@@ 3 Method
Figure 2: Overview of the proposed method. During model training, we first initialize known class embeddings with a pretrained CLIP text encoder [45]. The image encoder extracts a multi-scale feature map from the input. Then the RepVL-PAN [5] module uses multi-level cross-modal fusion to combine image and text features, forming the feature pyramids. The detection head predicts the class label based on image-text similarity and regresses the bounding box. The detection loss is used to update the known class embeddings. Concurrently, MSCAL modules are trained to maximize the consistency and similarity between class anchor and spatial locations at different scales, and output a multi-scale score map to indicate whether a feature is out of distribution (OOD) relative to a specified class. During the inference, the OOD map extracted by MSCAL is used to reduce known-unknown confusion. In addition, the pseudo unknown embedding used to discover unknown classes is constructed from the optimized known class embeddings and the generic “objectness” semantic concept.
@@
Open World Object Detection (OWOD) [19] aims to detect both known and unknown objects while continuously learning new classes. At an arbitrary stage $t$, we consider the known classes as $\mathcal{K}^{t}=\{1,...,N\}$, and unknown classes as $\mathcal{U}$. An OWOD model should be able to detect objects in $\mathcal{K}^{t}$ and $\mathcal{U}$, and extend known classes to $\mathcal{K}^{t+1}=\mathcal{K}^{t}\cup\{N+1,...,N+k\}$ when $k$ new classes are provided by human annotator. In this way, the object detector continuously discovers and learns new classes in the open world.
@@
Fig. 2 shows the general architecture of the proposed method. Following [5], we use text $T$ and image $I$ as inputs and match the text embeddings with image embeddings to predict class labels and bounding boxes of objects. Let $W_{\mathcal{K}}=\{w_{1},...,w_{N}\}$ denote the text embedding of $N$ known classes, which is initialized from class names encoded by the pre-trained CLIP [45] text encoder. $W_{\mathcal{K}}$ can be parameterized as model’s weight and optimized via Open World Embedding Learning (OWEL). During inference, a pseudo unknown embedding is constructed and appended to $W_{k}$, and the CLIP text encoder is disposable. The image encoder (DarkNet backbone inherited from Yolo v8 [18, 47]) extracts multi-scale features $C$ from the input image $I$. Then the multi-modal neck (RepVL-PAN [5]) uses multi-level cross-modal fusion to combine image and text features, forming the feature pyramids $P$. The detection head predicts bounding boxes and class labels by matching the cosine similarity of text embedding with each spatial location in $P$. Concurrently, MSCAL modules make dense predictions of out-of-distribution (OOD) scores and reject unknown bounding boxes in the detection head. Finally, the redundant predictions are filtered by Non-Maximum Suppression (NMS).
@@
As shown in Fig. 1, when an OVD model is given an unknown object, it will either misclassify it as a semantically similar text prompt, or ignore it if the semantic difference is large. For the former, we can model the distribution of known classes to reject out-of-distribution samples. For the latter, we need to construct a Pseudo Class Embedding $w_{\mathcal{U}}$ for unknown classes.
@@
Figure 4: Illustration of MSCAL module. For each layer in the feature pyramid, all spatial locations will be mapped to a new space and contrasted with class anchors. The design of the projector follows [63], which involves two 1 $\times$ 1 convolutional layers with ReLu and batch normalization. For a feature map $I\in\mathbb{R}^{H\times W\times C}$, the output will be $Z\in\mathbb{R}^{H\times W\times\frac{1}{2}C}$. The anchor is also parameterized as 1 $\times$ 1 convolutional layer. During inference, their inner product with the class anchor serves as the out-of-distribution score.
@@
$\mathcal{L}^{con}=\frac{1}{N}\mathcal{L}_{i}^{con}$ (4)
MSCAL contrasts class anchors with positive and negative samples across different scales and images. During the inference, the OOD score at spatial location $z$ is, $\mathcal{S}(z)=-\max_{i}\mu_{i}\cdot z$. The idea is that during training $\mu_{i}\cdot z$ will be maximized if $z$ belongs to class $i$, and minimized otherwise. Since an unknown object does not belong to any class, $\mathcal{S}(z)$ will be larger than that for known objects. We use this OOD score to identify unknown objects proposed by the detection head. The anchor $\mu$ can be parameterized as $1\times 1$ convolutional layer, which is computational efficient.
In addition, MSCAL is compatible with incremental learning. When the model encounters new objects and the training data for known classes is no longer available, we can simply add MSCAL modules for new classes and freeze modules for previously known classes to avoid catastrophic forgetting.
@@
Our method does not require exemplar replay to prevent catastrophic forgetting. We simply freeze the parameterized text embeddings and MSCAL modules for previous known classes and train new modules with currently known classes.
@@
For the OVD component of our model, we adopt the general architecture of Yolo World [5], which is pretrained on O365 [54], GoldG [20], CC3M [55]. Importantly, images overlapped with COCO dataset [27] has been removed from GoldG by default [5]. Our model is trained using the AdamW optimizer [31] with a learning rate of 1e-4 and a weight decay of 0.0125. The training batch size is set to 16, and the input images are rescaled to a resolution of 640×640. The weight parameter $\alpha$ is 0.4 for all tasks. All model training is conducted on a single Nvidia A100 GPU.
@@ 4 Experiments
Table 1: OWOD results on M-OWODB (top) and S-OWODB (bottom). Our method largely outperforms the SOTA methods in detecting both known mAP and unknown recall (U-Recall) objects on both benchmarks. The Base Model is the unmodified Yolo World detector [5] prompted with known class names and a hand-crafted generic object name (“object”). † uses a pretrained language model to learn the semantic topology of classes. * is based on a vision language model. Other results are directly taken from [58].
@@
Task IDs ($\rightarrow$) Task 1 Task 2 Task 3 Task 4 U-Recall mAP ($\uparrow$) U-Recall mAP ($\uparrow$) U-Recall mAP ($\uparrow$) mAP ($\uparrow$) Method ($\uparrow$) Current known ($\uparrow$) Previously known Current known Both ($\uparrow$) Previously known Current known Both Previously known Current known Both ORE [19] 4.9 56.0 2.9 52.7 26.0 39.4 3.9 38.2 12.7 29.7 29.6 12.4 25.3 OST [68] - 56.2 - 53.4 26.5 39.9 - 38.0 12.8 29.6 30.1 13.3 25.9 OW-DETR [14] 7.5 59.2 6.2 53.6 33.5 42.9 5.7 38.3 15.8 30.8 31.4 17.1 27.8 UC-OWOD [67] 2.4 50.7 3.4 33.1 30.5 31.8 8.7 28.8 16.3 24.6 25.6 15.9 23.2 ALLOW [35] 13.6 59.3 10.0 53.2 34.0 45.6 14.3 42.6 26.7 38.0 33.5 21.8 30.6 PROB [72] 19.4 59.5 17.4 55.7 32.2 44.0 19.6 43.0 22.2 36.0 35.7 18.9 31.5 CAT [33] 23.7 60.0 19.1 55.5 32.7 44.1 24.4 42.8 18.7 34.8 34.4 16.6 29.9 RandBox [64] 10.6 61.8 6.3 - - 45.3 7.8 - - 39.4 - - 35.4 EO-OWOD [58] 24.6 61.3 26.3 55.5 38.5 47.0 29.1 46.7 30.6 41.3 42.4 24.3 37.9 SKDF* [34] 39.0 56.8 36.7 52.3 28.3 40.3 36.1 36.9 16.4 30.1 31.0 14.7 26.9 Base Model 16.6 71.9 16.1 71.8 48.1 60.0 13.0 60.0 40.7 53.6 53.7 33.9 48.7 Ours 73.5 72.1 77.5 72.4 51.0 61.7 76.1 61.6 41.6 54.9 56.0 34.3 50.6 ORE [19] 1.5 61.4 3.9 56.5 26.1 40.6 3.6 38.7 23.7 33.7 33.6 26.3 31.8 OW-DETR [14] 5.7 71.5 6.2 62.8 27.5 43.8 6.9 45.2 24.9 38.5 38.2 28.1 33.1 PROB [72] 17.6 73.4 22.3 66.3 36.0 50.4 24.8 47.8 30.4 42.0 42.6 31.7 39.9 CAT [33] 24.0 74.2 23.0 67.6 35.5 50.7 24.6 51.2 32.6 45.0 45.4 35.1 42.8 EO-OWOD [58] 24.6 71.6 27.9 64.0 39.9 51.3 31.9 52.1 42.2 48.8 48.7 38.8 46.2 SKDF* [34] 60.9 69.4 60.0 63.8 26.9 44.4 58.6 46.2 28.0 40.1 41.8 29.6 38.7 Base Model 29.0 75.6 26.1 75.7 55.3 65.0 26.9 65.1 54.4 61.6 61.4 55.2 59.9 Ours 71.3 76.4 74.4 75.0 59.8 67.0 74.6 67.0 53.8 62.6 65.5 56.9 63.4
@@
Table 2: Uknown Object Confusion on M-OWODB. Wilderness impact (WI) and absolute open set error (A-OSE) reflect the negative impact of unknown objects on the accuracy of known classes. These metrics do not apply to task 4 since all classes are known.
@@
Task IDs ($\rightarrow$) Task 1 Task 2 Task 3 U-Recall WI A-OSE U-Recall WI A-OSE U-Recall WI A-OSE Method ($\uparrow$) ($\downarrow$) ($\downarrow$) ($\uparrow$) ($\downarrow$) ($\downarrow$) ($\uparrow$) ($\downarrow$) ($\downarrow$) ORE [19] 4.9 0.0621 10459 2.9 0.0282 10445 3.9 0.0211 7990 OST [68] - 0.0417 4889 - 0.0213 2546 - 0.0146 2120 OW-DETR [14] 7.5 0.0571 10240 6.2 0.0278 8441 5.7 0.0156 6803 PROB [72] 19.4 0.0569 5195 17.4 0.0344 6452 19.6 0.0151 2641 RandBox [64] 10.6 0.0240 4498 6.3 0.0078 1880 7.8 0.0054 1452 EO-OWOD [58] 24.6 0.0299 4148 26.3 0.0099 1791 29.1 0.0077 1345 Base Model 16.6 0.0311 9070 16.1 0.0147 7063 13.0 0.0086 5060 Ours 73.5 0.0175 1038 77.5 0.0047 529 76.1 0.0030 448

### v3

@@ Abstract
Traditional object detection methods operate under the closed-set assumption, where models can only detect a fixed number of objects predefined in the training set. Recent works on open vocabulary object detection (OVD) enable the detection of objects defined by an in-principle unbounded vocabulary, which reduces the cost of training models for specific tasks. However, OVD heavily relies on accurate prompts provided by an “oracle”, which limits their use in critical applications such as driving scene perception. OVD models tend to misclassify near-out-of-distribution (NOOD) objects that have similar features to known classes, and ignore far-out-of-distribution (FOOD) objects. To address these limitations, we propose a framework that enables OVD models to operate in open world settings, by identifying and incrementally learning previously unseen objects. To detect FOOD objects, we propose Open World Embedding Learning (OWEL) and introduce the concept of Pseudo Unknown Embedding which infers the location of unknown classes in a continuous semantic space based on the information of known classes. We also propose Multi-Scale Contrastive Anchor Learning (MSCAL), which enables the identification of misclassified unknown objects by promoting the intra-class consistency of object embeddings at different scales. The proposed method achieves state-of-the-art performance on standard open world object detection and autonomous driving benchmarks while maintaining its open vocabulary object detection capability.
@@ 1 Introduction
Recently, the rise of visual-language pre-training has spawned a new area: open vocabulary object detection (OVD). In principle, OVD can detect novel classes defined by an unbounded (open) vocabulary at inference. Naturally, it is able to incrementally learn new classes by adding new prompts. In the real world, however, the open vocabulary may not necessarily be complete and accurate. As shown in Fig. 1, an open vocabulary model such as Yolo World [5] is still a closed-set model at test time, because it is impractical for the expert user, referred to as “oracle”, to define all objects that the model might encounter in the real world. With a limited vocabulary the model tends to misclassify near-out-of-distribution (NOOD) objects that have similar semantics to known classes, and ignore far-out-of-distribution (FOOD) objects. In real-world applications such as autonomous driving, where countless object-types cannot be initially included in the text prompt, OVD will inevitably fail when presented with OOD objects.
To tackle this challenge, we propose a framework that enables OVD models to operate in open-world settings. To detect FOOD objects and incrementally learn the new classes, we propose Open World Embedding Learning (OWEL). OWEL optimizes parameterized class embeddings rather than fine-tuning the whole model to learn new classes, inherently avoiding catastrophic forgetting. We further introduce the novel concept of Pseudo Unknown Embedding, which constructs a text embedding to detect FOOD objects based on current known classes and generic objectness.
@@
Our method significantly surpasses state-of-the-art (SOTA) performance in U-Recall on the M-OWODB [19] and S-OWODB [14] benchmarks, while maintaining leading performance in other metrics. We further evaluate our method on a novel benchmark based on nuScenes [2], where it also achieves the best results. More importantly, our method preserves the zero-shot capability of the OVD model, as our implementation only optimizes text embeddings and additional MSCAL modules, while keeping the OVD model’s weights frozen. Therefore, our method provides a unified framework for both open vocabulary learning and open world learning. The contributions of this work are as follows:
  • We propose a framework that enables OVD models to operate in open world settings by identifying unknown objects and incrementally learning new classes, thereby unifying open vocabulary learning and open world learning within the same framework.
  • We propose a novel method, Open World Embedding Learning (OWEL), to enable the discovery and incremental learning of new classes without fine-tuning the whole model or requiring exemplars of previous tasks.
@@ 2 Related Works
Closed set object detection has been extensively studied over the past decade [12, 50, 11, 49, 30, 28, 61, 3]. To handle unknown objects and incrementally learn new classes, Joseph et al. [19] first proposed an open world object detector, which extends the Faster R-CNN [50] with contrastive clustering and energy-based unknown-object identification. Subsequent works [14, 36] used contextual information to improve unknown identification and knowledge transfer between known and unknown classes. To effectively detect unknown objects, Zohar et al. [75] proposed Probabilistic Objectness (PROB) to estimate the objectness of different proposals. Wang et al. [65] introduced random proposals in detector training to encourage the unknown discovery and reduce confusion between known and unknown classes. Ma et al. [34] proposed decoupling object localization and classification via cascade decoding. Sun et al. [59] further de-correlates objectness and class information by enforcing orthogonality. In some recent works [35, 37, 74], foundation models are employed in open-world learning.
@@
Out-of-distribution (OOD) detection is a concept closely related to open set learning, and the two terms are sometimes used interchangeably. OOD detection differs slightly from OSR on the choice of testing data [10, 52].
@@ 3 Method
Figure 2: Overview of the proposed method. During model training, we first initialize known class embeddings with a pretrained CLIP text encoder [46]. The image encoder extracts a multi-scale feature map from the input. Then the RepVL-PAN [5] uses multi-level cross-modal fusion to combine image and text features, forming the feature pyramids. The detection head predicts the class label based on image-text similarity and regresses the bounding box. The detection loss is used to update the known class embeddings. Concurrently, MSCAL modules are trained to maximize the similarity between class anchor and spatial locations at different scales, and output a multi-scale score map to indicate whether an embedding is out-of-distribution (OOD) relative to a specified class. During the inference, the OOD map extracted by MSCAL is used to reduce known-unknown confusion. In addition, the pseudo unknown embedding used to discover unknown classes is constructed from the optimized known class embeddings and the generic “objectness” semantic concept.
@@
Open World Object Detection (OWOD) [19] aims to detect both known and unknown objects while continuously learning new classes. At an arbitrary stage $t$, we consider the known classes as $\mathcal{K}^{t}=\{1,...,N\}$, and unknown classes as $\mathcal{U}$. An OWOD model should be able to detect objects in $\mathcal{K}^{t}$ and $\mathcal{U}$, and extend known classes to $\mathcal{K}^{t+1}=\mathcal{K}^{t}\cup\{N+1,...,N+k\}$ when $k$ new classes are incrementally learned. In this way, the object detector continuously discovers and learns new classes in the open world.
@@
Fig. 2 shows the general architecture of the proposed method. Following [5], we use text $T$ and image $I$ as inputs and match the text embeddings with image embeddings to predict class labels and bounding boxes of objects. Let $W_{\mathcal{K}}=\{w_{1},...,w_{N}\}$ denote the text embedding of $N$ known classes, which is initialized from class names encoded by the pre-trained CLIP [46] text encoder. $W_{\mathcal{K}}$ can be parameterized as model’s weight and optimized via Open World Embedding Learning (OWEL). During inference, a pseudo unknown embedding is constructed and appended to $W_{\mathcal{K}}$, and the CLIP text encoder is disposable. The image encoder (DarkNet backbone inherited from Yolo v8 [18, 48]) extracts multi-scale features $C$ from the input image $I$. Then the multi-modal neck (RepVL-PAN [5]) uses multi-level cross-modal fusion to combine image and text features, forming the feature pyramids $P$. The detection head predicts bounding boxes and class labels by matching the cosine similarity of text embedding with each spatial location in $P$. Concurrently, MSCAL modules make dense predictions of out-of-distribution (OOD) scores and reject unknown bounding boxes in the detection head. Finally, the redundant predictions are filtered by Non-Maximum Suppression (NMS).
@@
As shown in Fig. 1, when an OVD model is given an unknown object, it will either misclassify it as a semantically similar text prompt, or ignore it if the semantic difference is large. To avoid misclassification, we can model the distribution of known classes to reject out-of-distribution samples. To discover novel objects, we construct a Pseudo Class Embedding $w_{\mathcal{U}}$ representing unknown classes.
@@
Figure 4: Illustration of MSCAL module. For each layer in the feature pyramid, all spatial locations will be mapped to a new space and contrasted with class anchors. The design of the projector follows [64], which involves two 1 $\times$ 1 convolutional layers with ReLU and batch normalization. The anchor is also parameterized as 1 $\times$ 1 convolutional layer. During inference, their inner product with the class anchor serves as the OOD score.
@@
$\mathcal{L}^{con}=\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}_{i}^{con}$ (4)
MSCAL contrasts class anchors with positive and negative samples across different scales and images. During the inference, the OOD score at spatial location $z$ is, $\mathcal{S}(z)=-\max_{i}\mu_{i}\cdot z$. The idea is that during training $\mu_{i}\cdot z$ will be maximized if $z$ belongs to class $i$, and minimized otherwise. For an unknown object, $\mathcal{S}(z)$ will be larger than that for known objects. We use this OOD score to identify unknown objects proposed by the detection head. The anchor $\mu$ can be parameterized as $1\times 1$ convolutional layer, which is computational efficient.
MSCAL is also compatible with incremental learning. When the model encounters new objects, we can simply add MSCAL modules for new classes and freeze modules for previously known classes to avoid catastrophic forgetting.
@@
Our method does not require exemplar replay to prevent catastrophic forgetting. We simply freeze the parameterized text embeddings and MSCAL modules for previously known classes and train new modules with currently known classes.
@@
For the OVD component of our model, we adopt the general architecture of Yolo World [5], which is pretrained on O365 [55], GoldG [20], CC3M [56]. The backbone remains frozen in downstream tasks. Importantly, images overlapped with COCO dataset [27] has been removed from GoldG by default [5]. Our model is trained using the AdamW optimizer [32] with a learning rate of 1e-4 and a weight decay of 0.0125. The training batch size is set to 16, and the input images are rescaled to a resolution of 640×640. The weight parameter $\alpha$ is 0.4 for all tasks. All model training is conducted on a single Nvidia A100 GPU.
@@ 4 Experiments
Table 1: OWOD results on M-OWODB (top) and S-OWODB (bottom). Our method largely outperforms the SOTA methods in terms of both known mAP and unknown recall (U-Recall) on both benchmarks. The Base Model is the unmodified Yolo World detector [5] prompted with known class names and a hand-crafted generic object name (“object”). † uses a pretrained language model to learn the semantic topology of classes. * denotes models built on pretrained vision-language models. Other results are directly taken from [59].
@@
Task IDs ($\rightarrow$) Task 1 Task 2 Task 3 Task 4 U-Recall mAP ($\uparrow$) U-Recall mAP ($\uparrow$) U-Recall mAP ($\uparrow$) mAP ($\uparrow$) Method ($\uparrow$) Current known ($\uparrow$) Previously known Current known Both ($\uparrow$) Previously known Current known Both Previously known Current known Both ORE [19] 4.9 56.0 2.9 52.7 26.0 39.4 3.9 38.2 12.7 29.7 29.6 12.4 25.3 OST† [69] - 56.2 - 53.4 26.5 39.9 - 38.0 12.8 29.6 30.1 13.3 25.9 OW-DETR [14] 7.5 59.2 6.2 53.6 33.5 42.9 5.7 38.3 15.8 30.8 31.4 17.1 27.8 UC-OWOD [68] 2.4 50.7 3.4 33.1 30.5 31.8 8.7 28.8 16.3 24.6 25.6 15.9 23.2 ALLOW [36] 13.6 59.3 10.0 53.2 34.0 45.6 14.3 42.6 26.7 38.0 33.5 21.8 30.6 PROB [75] 19.4 59.5 17.4 55.7 32.2 44.0 19.6 43.0 22.2 36.0 35.7 18.9 31.5 CAT [34] 23.7 60.0 19.1 55.5 32.7 44.1 24.4 42.8 18.7 34.8 34.4 16.6 29.9 RandBox [65] 10.6 61.8 6.3 - - 45.3 7.8 - - 39.4 - - 35.4 EO-OWOD [59] 24.6 61.3 26.3 55.5 38.5 47.0 29.1 46.7 30.6 41.3 42.4 24.3 37.9 MAVL* [37] 50.1 64.0 49.5 61.6 30.8 46.2 50.9 43.8 22.7 36.8 36.2 20.6 32.3 SKDF* [35] 39.0 56.8 36.7 52.3 28.3 40.3 36.1 36.9 16.4 30.1 31.0 14.7 26.9 Base Model 16.6 71.9 16.1 71.8 48.1 60.0 13.0 60.0 40.7 53.6 53.7 33.9 48.7 Ours 73.5 72.1 77.5 72.4 51.0 61.7 76.1 61.6 41.6 54.9 56.0 34.3 50.6 ORE [19] 1.5 61.4 3.9 56.5 26.1 40.6 3.6 38.7 23.7 33.7 33.6 26.3 31.8 OW-DETR [14] 5.7 71.5 6.2 62.8 27.5 43.8 6.9 45.2 24.9 38.5 38.2 28.1 33.1 PROB [75] 17.6 73.4 22.3 66.3 36.0 50.4 24.8 47.8 30.4 42.0 42.6 31.7 39.9 CAT [34] 24.0 74.2 23.0 67.6 35.5 50.7 24.6 51.2 32.6 45.0 45.4 35.1 42.8 EO-OWOD [59] 24.6 71.6 27.9 64.0 39.9 51.3 31.9 52.1 42.2 48.8 48.7 38.8 46.2 SKDF* [35] 60.9 69.4 60.0 63.8 26.9 44.4 58.6 46.2 28.0 40.1 41.8 29.6 38.7 Base Model 29.0 75.6 26.1 75.7 55.3 65.0 26.9 65.1 54.4 61.6 61.4 55.2 59.9 Ours 71.3 76.4 74.4 75.0 59.8 67.0 74.6 67.0 53.8 62.6 65.5 56.9 63.4
@@
Table 2: Unknown Object Confusion on M-OWODB. Wilderness impact (WI) and absolute open set error (A-OSE) reflect the negative impact of unknown objects on the accuracy of known classes. These metrics do not apply to task 4 since all classes are known.
@@
Task IDs ($\rightarrow$) Task 1 Task 2 Task 3 U-Recall WI A-OSE U-Recall WI A-OSE U-Recall WI A-OSE Method ($\uparrow$) ($\downarrow$) ($\downarrow$) ($\uparrow$) ($\downarrow$) ($\downarrow$) ($\uparrow$) ($\downarrow$) ($\downarrow$) ORE [19] 4.9 0.0621 10459 2.9 0.0282 10445 3.9 0.0211 7990 OST† [69] - 0.0417 4889 - 0.0213 2546 - 0.0146 2120 OW-DETR [14] 7.5 0.0571 10240 6.2 0.0278 8441 5.7 0.0156 6803 PROB [75] 19.4 0.0569 5195 17.4 0.0344 6452 19.6 0.0151 2641 RandBox [65] 10.6 0.0240 4498 6.3 0.0078 1880 7.8 0.0054 1452 EO-OWOD [59] 24.6 0.0299 4148 26.3 0.0099 1791 29.1 0.0077 1345 Base Model 16.6 0.0311 9070 16.1 0.0147 7063 13.0 0.0086 5060 Ours 73.5 0.0175 1038 77.5 0.0047 529 76.1 0.0030 448